# ShiftAddViT: Mixture of Multiplication Primitives Towards Efficient Vision Transformers

### v4

@@ Appendix A More Speedup Comparisons of Shift and Add
Apart from the speedup analysis when the batch size is one, we also conducted GPU speedup tests on our customized Shift and Add kernels when using larger batch sizes, such as 32, as depicted in Fig. 8 and 8, respectively. The results demonstrate that both our customized kernels consistently outperform the PyTorch and TVM baselines in terms of speed. For instance, our MatAdds achieve an average speedup of 2.80$\times$/1.65$\times$ compared to PyTorch and TVM MatMuls, respectively. Similarly, our MatShifts achieve an average speedup of 3.38$\times$/1.21$\times$ compared to TVM FakeShifts elhoushi2021deepshift and TVM MatMuls, respectively. It is important to note that in these comparisons, we ensured fair conditions by using the default einsum operator for PyTorch MatMuls, employing floating-point multiplication with power-of-twos for FakeShift, and implementing the MatAdds and MatShifts ourselves using TVM chen2018tvm . We compared these customized operators with their multiplication counterparts in the same setup to ensure a fair evaluation. The observed speedup of MatAdds is due to the replacement of multiplication-and-accumulation with accumulation only. On the other hand, the speedup of MatShifts is primarily attributed to the reduction in bit precision (INT32 and INT8 for inputs and shift signs/weights, respectively), resulting in reduced data movement instead of computations. These computational improvements are almost entirely hidden behind data movements.
Figure 7: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs PVT_ICCV .
Figure 7: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs PVT_ICCV .
@@ Appendix C More Results on 3D Tasks
We have extended our methods to 3D NVS tasks and have built ShiftAddViT on top of Transformer-based GNT models GNT . In the main paper, we tested and reported the averaged PSNR, SSIM, LPIPS, as well as latency and energy of both ShiftAddViT and baselines, i.e., the vanilla NeRF mildenhall2021nerf and GNT GNT , on the LLFF dataset across eight scenes. Here, we further supply the full results of eight scenes. As shown in Tab. 8, 9, 10, our ShiftAddViT consistently leads to better accuracy-efficiency tradeoffs in terms of the three metrics mentioned above, achieving comparable or even better generation quality ($\uparrow$0.55$\downarrow$0.19 averaged PSNR across eight scenes), as compared to NeRF and GNT baselines, respectively. In particular, ShiftAddViT achieves better generation quality (up to $\uparrow$0.88 PSNR) than GNT on some representation scenes, e.g., Orchid and Flower. In all tables, the first, second, and third ranking performances are noted with corresponding colors.
@@
Figure 10: Qualitative comparisons between (a) NeRF mildenhall2021nerf , (b) GNT GNT , (c) ShiftAddViT (Ours), and (d) ground truth when tested on the LLFF dataset for single-scene rendering.
Table 8: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT GNT , and report averaged PSNR ($\uparrow$) on the LLFF dataset mildenhall2019local across eight scenes.
  Methods Add Shift MoE PSNR (normal-$\uparrow$) on LLFF Dataset
@@
NeRF mildenhall2021nerf - - - 32.70 25.17
20.92
31.16
20.36
27.40
26.80
27.45
GNT GNT - - - 32.96
24.31
22.57 32.28
20.67
27.32
28.15 29.62
ShiftAddViT
31.68
24.63
22.34
31.79
20.74
28.02
26.99
28.62
✓(Both)
31.64
24.63
22.34
31.73
20.78 28.05
27.03
28.61
✓(Attn) ✓(MLPs)
31.85
24.65 22.37 31.85
20.73
28.20 27.08 28.62
✓(Both) 32.12
24.35
22.42 32.15 20.84 28.14 27.32 29.09
Table 9: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT GNT , and report averaged SSIM ($\uparrow$) on the LLFF dataset mildenhall2019local across eight scenes.
  Methods Add Shift MoE SSIM (normal-$\uparrow$) on LLFF Dataset
@@
NeRF mildenhall2021nerf - - -
0.948
0.792
0.690
0.881
0.641
0.827
0.880
0.828
GNT GNT - - - 0.963 0.846 0.852 0.934 0.752 0.893 0.936 0.935
ShiftAddViT
0.949
0.839
0.838
0.922
0.730
0.891
0.916
0.911
✓(Both)
0.948
0.839
0.839
0.923
0.730
0.892
0.917
0.911
@@
✓(Both) 0.955
0.836
0.841 0.930 0.746 0.896 0.923 0.922
Table 10: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT GNT , and report averaged LPIPS ($\downarrow$) on the LLFF dataset mildenhall2019local across eight scenes.
  Methods Add Shift MoE LPIPS (normal-$\downarrow$) on LLFF Dataset
@@
NeRF mildenhall2021nerf - - -
0.178
0.280
0.316
0.171
0.321
0.219
0.249
0.268
GNT GNT - - - 0.060 0.116 0.109 0.061 0.153
0.092
0.080 0.076
ShiftAddViT
0.087
0.141
0.129
0.080
0.182
0.089
0.107
0.110
✓(Both)
0.088
0.142
0.128
0.080
0.182
0.088
0.106
0.110
@@
In addition, we supply more qualitative rendering visualization on three scenes, Horns, Orchids, and Flowers, as shown in Fig. 10. We see that our ShiftAddViT generates more clear images than vanilla NeRF mildenhall2021nerf and comparable or even better details than GNT GNT . Quantitative and qualitative results consistently show that our ShiftAddViT does not hurt rendering quality when improving rendering speeds and energy efficiency.
@@ Appendix D More Results on NLP Tasks
We also test our proposed optimization of MHSA and MLPs to NLP tasks. In particular, we apply our methods to Transformer models and Long Range Arena (LRA) benchmarks consisting of sequences ranging from 1K to 16K tokens tay2020long . The results are shown in Tab. 11. We see that our proposed ShiftAdd Transformers consistently show superior performance in terms of both model accuracy (+0.45% $\sim$ +5.79%) and efficiency (1.5$\times$ $\sim$ 11.5$\times$ latency reductions and 2.2$\times$ $\sim$ 16.4$\times$ energy reductions on an Eyeriss-like accelerator) as compared to original Transformer and other linear attention baselines, which means that our shift and add reparameterization and load-balanced MoE ideas are generally applicable to Transformer models and are agnostic to domains and tasks.
@@ Appendix E More Experiment Settings
For the classification task, we first replace the original attention layers in PVTv1 PVT_ICCV and PVTv2 wang2022pvt with the standard MSA, and then finetune the MSA-based variants on top of the ImageNet-$1$k-pretrained weights for $100$ epochs with a learning rate (lr) of $5\times 10^{-5}$ following Ecoformer liu2022ecoformer . After that, we reparameterize the MSA-based models through a two-stage finetuning process: (1) convert MSA to linear attention you2023castling and reparameterize MatMuls with add layers via $100$ epochs of finetuning with a base lr of $1\times 10^{-5}$, and (2) reparameterize MLPs or linear layers with shift or MoE layers through another $100$ epochs of finetuning with a base lr of $1\times 10^{-5}$. All models are trained using eight RTX A5000 GPUs with a total batch size of $256$, and we use the AdamW optimizer Loshchilov2017DecoupledWD with a cosine decay lr scheduler. All other hyperparameters are the same as those in Ecoformer liu2022ecoformer and PVTv2 wang2022pvt . For the NVS task, we apply reparameterization on top of the pretrained models from GNT GNT using a similar two-stage fine-tuning process, except that we do not convert MSA to linear attention for maintaining accuracy. Specifically, we first (1) reparameterize MatMuls with add layers and then (2) reparameterize MLPs or linear layers with shift or MoE layers. Both stages are finetuned for $140$K steps with a base lr of $5\times 10^{-4}$, and we sample $2048$ rays with $192$ coarse points sampled per ray in each iteration. All other hyperparameters are the same as those in GNT GNT , including the use of the Adam optimizer with an exponential decay lr scheduler.
Scaling Factor Clarification. For shift layers, we reparameterize shifts following DeepShift-PS elhoushi2021deepshift and do not use a scaling factor. A straight-through estimator (STE) yin2019understanding is adopted to make the shift layer differentiable. For ShiftAddViT with KSH liu2022ecoformer , there is no scaling factor needed as a set of hash functions is applied to convert Q/K to binary codes. For ShiftAddViT with Quant. hubara2017quantized , we leverage layer-wise Quant. for both Q & K, the scaling factor can be multiplied after add operations. It can be efficiently implemented following jacob2018quantization .
@@ Appendix F More Ablation Studies
The Advantage of Using Shift. Replacing Multiplication-based MLPs with shift layers significantly reduces latency, energy, and chip area costs. Our shift kernel offers an average speedup of 2.35$\times$/3.07$\times$/1.16$\times$ compared to PyTorch FakeShift, TVM FakeShift, and TVM MatMuls, respectively, as shown in Fig. 3. The seemingly minor latency improvements in Tab. 4 and 6 are due to full optimization of the compiled model as a whole (e.g., 6.34ms normal-$\rightarrow$ 1ms for PVTv2-B0) on GPUs with sufficient chip area. Most gains are concealed by data movements and system-level schedules. Adopting shift layers substantially lowers energy and chip area usage as validated in Tab. 1. Under the same chip areas, latency savings are more pertinent, e.g., PVTv2-B0 w/ shift or MoE achieve 3.9 $\sim$ 5.7$\times$ and 1.6 $\sim$ 1.8$\times$ speedups, respectively, as summarized in Tab. 13.

### v6

@@ Appendix A More Speedup Comparisons of Shift and Add
Apart from the speedup analysis when the batch size is one, we also conducted GPU speedup tests on our customized Shift and Add kernels when using larger batch sizes, such as 32, as depicted in Fig. 8 and 8, respectively. The results demonstrate that both our customized kernels consistently outperform the PyTorch and TVM baselines in terms of speed. For instance, our MatAdds achieve an average speedup of 2.80$\times$/1.65$\times$ compared to PyTorch and TVM MatMuls, respectively. Similarly, our MatShifts achieve an average speedup of 3.38$\times$/1.21$\times$ compared to TVM FakeShifts [17] and TVM MatMuls, respectively. It is important to note that in these comparisons, we ensured fair conditions by using the default einsum operator for PyTorch MatMuls, employing floating-point multiplication with power-of-twos for FakeShift, and implementing the MatAdds and MatShifts ourselves using TVM [10]. We compared these customized operators with their multiplication counterparts in the same setup to ensure a fair evaluation. The observed speedup of MatAdds is due to the replacement of multiplication-and-accumulation with accumulation only. On the other hand, the speedup of MatShifts is primarily attributed to the reduction in bit precision (INT32 and INT8 for inputs and shift signs/weights, respectively), resulting in reduced data movement instead of computations. These computational improvements are almost entirely hidden behind data movements.
Figure 7: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs [60].
@@ Appendix C More Results on 3D Tasks
We have extended our methods to 3D NVS tasks and have built ShiftAddViT on top of Transformer-based GNT models [53]. In the main paper, we tested and reported the averaged PSNR, SSIM, LPIPS, as well as latency and energy of both ShiftAddViT and baselines, i.e., the vanilla NeRF [41] and GNT [53], on the LLFF dataset across eight scenes. Here, we further supply the full results of eight scenes. As shown in Tab. 8, 9, 10, our ShiftAddViT consistently leads to better accuracy-efficiency tradeoffs in terms of the three metrics mentioned above, achieving comparable or even better generation quality ($\uparrow$0.55$\downarrow$0.19 averaged PSNR across eight scenes), as compared to NeRF and GNT baselines, respectively. In particular, ShiftAddViT achieves better generation quality (up to $\uparrow$0.88 PSNR) than GNT on some representation scenes, e.g., Orchid and Flower. In all tables, the first, second, and third ranking performances are noted with corresponding colors.
@@
Figure 10: Qualitative comparisons between (a) NeRF [41], (b) GNT [53], (c) ShiftAddViT (Ours), and (d) ground truth when tested on the LLFF dataset for single-scene rendering.
Table 8: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT [53], and report averaged PSNR ($\uparrow$) on the LLFF dataset [40] across eight scenes.
  Methods Add Shift MoE PSNR ($\uparrow$) on LLFF Dataset
@@
NeRF [41] - - - 32.70 25.17 20.92 31.16 20.36 27.40 26.80 27.45
GNT [53] - - - 32.96 24.31 22.57 32.28 20.67 27.32 28.15 29.62
ShiftAddViT 31.68 24.63 22.34 31.79 20.74 28.02 26.99 28.62
✓(Both) 31.64 24.63 22.34 31.73 20.78 28.05 27.03 28.61
✓(Attn) ✓(MLPs) 31.85 24.65 22.37 31.85 20.73 28.20 27.08 28.62
✓(Both) 32.12 24.35 22.42 32.15 20.84 28.14 27.32 29.09
Table 9: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT [53], and report averaged SSIM ($\uparrow$) on the LLFF dataset [40] across eight scenes.
  Methods Add Shift MoE SSIM ($\uparrow$) on LLFF Dataset
@@
NeRF [41] - - - 0.948 0.792 0.690 0.881 0.641 0.827 0.880 0.828
GNT [53] - - - 0.963 0.846 0.852 0.934 0.752 0.893 0.936 0.935
ShiftAddViT 0.949 0.839 0.838 0.922 0.730 0.891 0.916 0.911
✓(Both) 0.948 0.839 0.839 0.923 0.730 0.892 0.917 0.911
@@
✓(Both) 0.955 0.836 0.841 0.930 0.746 0.896 0.923 0.922
Table 10: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of the SOTA Transformer-based NeRF model, GNT [53], and report averaged LPIPS ($\downarrow$) on the LLFF dataset [40] across eight scenes.
  Methods Add Shift MoE LPIPS ($\downarrow$) on LLFF Dataset
@@
NeRF [41] - - - 0.178 0.280 0.316 0.171 0.321 0.219 0.249 0.268
GNT [53] - - - 0.060 0.116 0.109 0.061 0.153 0.092 0.080 0.076
ShiftAddViT 0.087 0.141 0.129 0.080 0.182 0.089 0.107 0.110
✓(Both) 0.088 0.142 0.128 0.080 0.182 0.088 0.106 0.110
@@
In addition, we supply more qualitative rendering visualization on three scenes, Horns, Orchids, and Flowers, as shown in Fig. 10. We see that our ShiftAddViT generates more clear images than vanilla NeRF [41] and comparable or even better details than GNT [53]. Quantitative and qualitative results consistently show that our ShiftAddViT does not hurt rendering quality when improving rendering speeds and energy efficiency.
@@ Appendix D More Results on NLP Tasks
We also test our proposed optimization of MHSA and MLPs to NLP tasks. In particular, we apply our methods to Transformer models and Long Range Arena (LRA) benchmarks consisting of sequences ranging from 1K to 16K tokens [54]. The results are shown in Tab. 11. We see that our proposed ShiftAdd Transformers consistently show superior performance in terms of both model accuracy (+0.45% $\sim$ +5.79%) and efficiency (1.5$\times$ $\sim$ 11.5$\times$ latency reductions and 2.2$\times$ $\sim$ 16.4$\times$ energy reductions on an Eyeriss-like accelerator) as compared to original Transformer and other linear attention baselines, which means that our shift and add reparameterization and load-balanced MoE ideas are generally applicable to Transformer models and are agnostic to domains and tasks.
@@ Appendix E More Experiment Settings
For the classification task, we first replace the original attention layers in PVTv1 [60] and PVTv2 [61] with the standard MSA, and then finetune the MSA-based variants on top of the ImageNet-$1$k-pretrained weights for $100$ epochs with a learning rate (lr) of $5\times 10^{-5}$ following Ecoformer [34]. After that, we reparameterize the MSA-based models through a two-stage finetuning process: (1) convert MSA to linear attention [73] and reparameterize MatMuls with add layers via $100$ epochs of finetuning with a base lr of $1\times 10^{-5}$, and (2) reparameterize MLPs or linear layers with shift or MoE layers through another $100$ epochs of finetuning with a base lr of $1\times 10^{-5}$. All models are trained using eight RTX A5000 GPUs with a total batch size of $256$, and we use the AdamW optimizer [37] with a cosine decay lr scheduler. All other hyperparameters are the same as those in Ecoformer [34] and PVTv2 [61]. For the NVS task, we apply reparameterization on top of the pretrained models from GNT [53] using a similar two-stage fine-tuning process, except that we do not convert MSA to linear attention for maintaining accuracy. Specifically, we first (1) reparameterize MatMuls with add layers and then (2) reparameterize MLPs or linear layers with shift or MoE layers. Both stages are finetuned for $140$K steps with a base lr of $5\times 10^{-4}$, and we sample $2048$ rays with $192$ coarse points sampled per ray in each iteration. All other hyperparameters are the same as those in GNT [53], including the use of the Adam optimizer with an exponential decay lr scheduler.
Scaling Factor Clarification. For shift layers, we reparameterize shifts following DeepShift-PS [17] and do not use a scaling factor. A straight-through estimator (STE) [69] is adopted to make the shift layer differentiable. For ShiftAddViT with KSH [34], there is no scaling factor needed as a set of hash functions is applied to convert Q/K to binary codes. For ShiftAddViT with Quant. [27], we leverage layer-wise Quant. for both Q & K, the scaling factor can be multiplied after add operations. It can be efficiently implemented following [28].
@@ Appendix F More Ablation Studies
The Advantage of Using Shift. Replacing Multiplication-based MLPs with shift layers significantly reduces latency, energy, and chip area costs. Our shift kernel offers an average speedup of 2.35$\times$/3.07$\times$/1.16$\times$ compared to PyTorch FakeShift, TVM FakeShift, and TVM MatMuls, respectively, as shown in Fig. 3. The seemingly minor latency improvements in Tab. 4 and 6 are due to full optimization of the compiled model as a whole (e.g., 6.34ms $\rightarrow$ 1ms for PVTv2-B0) on GPUs with sufficient chip area. Most gains are concealed by data movements and system-level schedules. Adopting shift layers substantially lowers energy and chip area usage as validated in Tab. 1. Under the same chip areas, latency savings are more pertinent, e.g., PVTv2-B0 w/ shift or MoE achieve 3.9 $\sim$ 5.7$\times$ and 1.6 $\sim$ 1.8$\times$ speedups, respectively, as summarized in Tab. 13.